In a double-slit experiment, the seventh-order bright fringe is observed at an angle of 1.66°. If the slit separation is 0.11 mm, then what is the wavelength of the light?
Answer #1

Here,

let the wavelength is L

as d * sin(theta) = m * wavelength

0.11 *10^-3 * sin(1.66 degree) = 7 * wavelength

solving for wavelength

wavelength = 4.55 *10-7 m = 455 nm

the wavelength of light used is 455 nm